Jasmine is three rimes as old as Jasper. In 12 years, Jasmine will be twice as old as Jasper. Find the present age of Jasper?
Answer by mananth(16949)

Jasper age =x
Jasmine age = 3x
their ages after 12 years
jasper = x+12
Jasmine = 3x+12
3x+12 =2(x+12)
3x+12=2x+24
3x-2x=24-12
x=12 jasper's age